OpenStudy (anonymous):

If there are 1,200 ounces calories in 8 ounces of hot fudge,how many calories are in 3 ounces of hot fudge?

OpenStudy (aripotta):

set up a proportion:\[\frac{ calories }{ ounces }\frac{ 1200 }{ 8 }=\frac{ x }{ 3 }\]then cross-multiply and solve for x